Initially, 1.0 mol of NO(g) and 1 mol of Cl₂(g) were added to a 1 L container. As a result of the reaction the equilibrium concentration of NOCl(g) became 0.96 M. Using the ICE table methodology determine the value of the equilibrium constant KC for this reaction. . 2NO(g) + Cl₂(g) ⇋2NOCl(g)

Your answer



1
Expert's answer
2021-03-11T08:46:57-0500

2NO(g) + Cl2(g) "\leftrightarrow" 2NOCl

1mol/L____1mol/L____0mol/L

-2x______-x_________+2x


[NOCl]eq = 2x = 0.96M


x = 0.48 M


Kc = "\frac{[NOCl]^2}{[NO]^2[Cl]}" = "\frac{[0.48]^2}{[1-2x]^2[1-x]}"


Kc = "\frac{0.2304}{[1-0.96]^2[1-0.48]}" = "\frac{0.2304}{[0.04]^2[0.52]}"


Kc = "\frac{0.2304}{(0.0016)×0.52}" = "\frac{0.2304}{0.000832}" = 276.92